### v1

# Unifying Appearance Codes and Bilateral Grids for Driving Scene Gaussian Splatting

Abstract
Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions. However, in real-world scenarios, it is challenging to guarantee perfect photometric consistency in acquired images. Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image. Recently, the bilateral grid was introduced to perform pixel-wise color mapping, but it is difficult to optimize and constrain effectively. In this paper, we propose a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids. We demonstrate that this approach significantly improves geometric accuracy in dynamic, decoupled autonomous driving scene reconstruction, outperforming both appearance codes and bilateral grids. This is crucial for autonomous driving, where accurate geometry is important for obstacle avoidance and control. Our method shows strong results across four datasets: Waymo, NuScenes, Argoverse, and PandaSet. We further demonstrate that the improvement in geometry is driven by the multi-scale bilateral grid, which effectively reduces floaters caused by photometric inconsistency. Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
Figure 1: (a) Appearance codes rely on global affine transformations but have limited modeling capability. (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize. (c) The proposed multi-scale bilateral grid unifies appearance codes and bilateral grids, enabling patch-wise transformations.
1 Introduction
Neural rendering techniques, such as NeRFs  mildenhall2021nerf ; wang2021neus ; yu2022monosdf ; liu2024rip ; yuan2024slimmerf and Gaussian Splatting (GS) kerbl20233d ; chen2024pgsr ; cheng2024gaussianpro ; zhang2024drone ; song2024sa , have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency. However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably martin2021nerf ; zhang2024gaussian ; kulhanek2024wildgaussians ; wang2025semantic . In autonomous driving applications yan2024renderworld ; wu2023mars ; tonderski2024neurad ; chen2024omnire ; li2024uniscene ; jin2024tod3cap ; tian2023unsupervised , these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time. As illustrated in the left half of Teaser. 1, we observe that these variations—whether temporal (across different time steps) or inter-camera (across different viewpoints)—can introduce significant discrepancies in image appearance. These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
To address this issue, appearance codes tonderski2024neurad ; martin2021nerf ; chen2024omnire have been introduced to encode per-image information and assist in correcting the photometric discrepancies. These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes. For example, in dynamic scenes, such as those encountered in autonomous driving environments, lighting and viewpoints can change across different camera angles and time frames. As shown in the right top of Teaser. 1, a global affine transformation based on a single appearance code may not be sufficient to handle such variations (In Teaser. 1 NuScenes (a), the fence is blurred).
Recent advancements wang2024bilateral ; gharbi2017deep have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments. However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser. 1 (b), bilateral grids fail to effectively optimize the reconstruction of large, complex scenes.). In this work, we propose a novel solution—a multi-scale bilateral grid—integrating the strengths of appearance codes and bilateral grids. As shown in Teaser. 1, this new approach allows for patch-wise transformations. Interestingly, in extreme cases, the multi-scale bilateral grid naturally converges to either the bilateral grid or appearance code, depending on the scale. At the finest scale, the multi-scale grid behaves like a traditional bilateral grid, performing pixel-wise transformations to adjust for local variations. In contrast, at the coarsest scale, it effectively reverts to a more global transformation, resembling the behavior of appearance codes. This flexible, scale-dependent approach offers the best of both worlds, providing localized fine-tuning where necessary, while maintaining global consistency when appropriate.
We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning. This is evidenced by our evaluation on multiple widely used datasets, including Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset . The corresponding Chamfer Distance (CD) values, displayed in the Teaser. 1, highlight a marked reduction in error and improvement in reconstruction quality. Additionally, our method reduces photometric inconsistency, mitigating the appearance of floaters and enhancing the overall realism of the scene.
The contributions of this paper are as follows: First, we introduce a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids, transitioning to either of the two paradigms in extreme cases, thus enhancing both modeling capability and optimization efficiency. Second, we show that by addressing photometric inconsistencies, it improves the geometric accuracy of dynamic, decoupled driving scene reconstructions. Third, we provide extensive benchmarking across four widely used datasets—Waymo, NuScenes, Argoverse, and PandaSet—where our method outperforms previous approaches, showcasing notable improvements in both qualitative and quantitative results.
Figure 2: Overview of our method. We unify appearance codes with multi-scale bilateral grids. Initially, a coarse rendering is obtained from a Gaussian scene graph. This rendered image is then processed by our multi-scale bilateral grids to perform detailed per-pixel color modeling, guided by a luminance-based map through slice and fusion operations.
2 Related Works
2.1 Bilateral Grids and Appearance Codes
Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving. The bilateral grids, first introduced for real-time edge-aware image processing chen2007real , enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques gharbi2017deep ; marki2016bilateral ; chen2016bilateral ; xu2021bilateral ; zheng2021ultra . Extensions like deep bilateral learning gharbi2017deep further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices. Similarly, bilateral-guided upsampling chen2016bilateral and cost volume refinement xu2021bilateral have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations. In the context of neural rendering, methods such as NeRF tancik2022block ; jang2021codenerf ; martin2021nerf and Gaussian Splatting (GS) zhang2024gaussian ; kulhanek2024wildgaussians have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders martin2021nerf ; zhang2024gaussian remains a challenge. Techniques like bilateral guided radiance field processing wang2024bilateral address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space. WildGaussians kulhanek2024wildgaussians integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes. Cross-Ray NeRF yang2023cross employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections. Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability. For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations durand2002fast . Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively paris2006fast . To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
2.2 Autonomous Driving Simulation
Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios wu2023mars ; okae2021robust ; yan2024street ; lindstrom2024nerfs ; tao2024alignmif ; li2024uniscene . Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation wu2023mars ; li2024hierarchical ; zhou2024drivinggaussian ; tonderski2024neurad ; yan2024street . For instance, MARS wu2023mars employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian yan2024street achieves real-time urban scene rendering using explicit 3D Gaussian representations. DrivingGaussian zhou2024drivinggaussian enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency. NeuRAD tonderski2024neurad integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis. Holistic scene understanding and editing have also seen progress. HUGS zhou2024hugs combines static and dynamic 3D Gaussians for real-time semantic parsing. ChatSim wei2024editable enables language-driven scene editing with external asset integration. LiDAR data integration has advanced through methods like LiDAR-NeRF tao2024lidar , which uses structural regularization for low-texture regions. HO-Gaussian li2024ho , which merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion (SFM) dependencies.
Despite progress, gaps persist between simulated and real-world data lindstrom2024nerfs ; tao2024alignmif . AlignMiF tao2024alignmif addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS nguyen2024rodus decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts. Innovations in large-scale generation include InfiniCube lu2024infinicube , which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire chen2024omnire , which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations. Furthermore, recent works like GaussianPro cheng2024gaussianpro and SplatAD hess2024splatad explore to refine 3D GS for real-time rendering of dynamic scenes. GaussianPro cheng2024gaussianpro introduces progressive propagation for texture-less surfaces, and SplatAD hess2024splatad models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity). In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
3 Methodology
3.1 Problem Formulation
We aim to reconstruct a 3D scene representation $\mathbb{G}$ from multi-view images and LiDAR depth maps, commonly available in autonomous driving datasets. Given a set of images $\{I_{c,t}\}$ and corresponding depths $\{D_{c,t}\}$ captured by $N$ cameras over $T$ time steps, we formulate our objective as minimizing the discrepancy between rendered and observed images and depths:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|I_{c,t}^{r}-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~{}~{},$ (1)
where $I_{c,t}^{r}$ and $D_{c,t}^{r}$ are the rendered appearance and geometry of the scene $\mathbb{G}$ from camera $c$ at time $t$, and $\lambda_{D}$ is a weight balancing the depth term.
However, dynamic driving scenes and varying camera properties cause inconsistent appearances, leading to geometric and texture artifacts. To address inconsistency, we decompose each image $I_{c,t}$ into consistent $\mathcal{C}_{c,t}$ (e.g., intrinsic scene colors and constant sensor adjustments like normalization exposure), and non-consistent $\mathcal{N}_{c,t}$ (e.g., components-varying lighting conditions, different camera settings and diverse image ISP effects):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{N}_{c,t}~{}~{},$ (2)
We model $\mathcal{N}_{c,t}$ as a non-linear transformation $\mathcal{F}(\mathcal{C}_{c,t})$. Substituting into Eq. (2):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})~{}~{},$ (3)
Let $\mathcal{E}(\mathcal{C}_{c,t})=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})$, we can reformulate the optimization objective in Eq. (1) as:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|\mathcal{E}(\mathcal{C}_{c,t}^{r})-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~{}~{},$ (4)
This reformulation links photometric and geometric consistency optimization within a joint objective, guiding $\mathcal{C}_{c,t}^{r}$ adjustments by geometric constraints ($D_{c,t}^{r}$). This mitigates texture-geometry ambiguities (e.g., shadows on road surfaces misinterpretable as geometry changes). $\mathcal{F}(\mathcal{C}_{c,t})$ models transient appearance, while $\mathcal{C}_{c,t}^{r}$ enforces consistent scene properties. Experiments (Sec. 4, Tab. 1) show this joint optimization reduces both photometric error and geometric drift in comparison to baselines.
3.2 Gaussian Splatting for Autonomous Driving Environments
Following prior methodologies chen2024omnire ; tonderski2024neurad , we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models. Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations. We derive the transformation matrixes from existing object detection pipelines or ground-truth annotations.
For the sky, we use an environment map to model sky color based on viewing direction, while static background is represented as a set of semi-transparent 3D Gaussians. Each Gaussian is characterized by a learnable opacity parameter $o\in(0,1)$, a mean position $\mu\in\mathbb{R}^{3}$, and an anisotropic covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ which is parameterized by a scale vector $S\in\mathbb{R}^{3}$ and a rotation quaternion $q\in\mathbb{R}^{4}$. Additionally, Spherical harmonics coefficients $c\in\mathbb{R}^{F}$ are used to model appearance.
For the moving object models, we further distinguish between non-deformable (e.g., vehicles) and deformable objects (e.g., pedestrians). Non-deformable objects $\{\mathbb{G}_{i}^{N}|i\in\{1,\dots,n_{n}\}\}$ are optimized in their local coordinate space and transformed into the global world space via their pose $T$:
$\mathbb{G}_{i}^{N}=T\otimes\hat{\mathbb{G}}_{i}^{N}~{}~{},$ (5)
For deformable objects $\{\mathbb{G}_{i}^{D}|i\in\{1,\dots,n_{d}\}\}$, we employ a deformation network $\mathcal{F}_{\Phi}$ (parameterized by $\Phi$) to adapt the Gaussian representation based on latent variable $e$ and time $t$:
$\mathbb{G}_{i}^{D}=T\otimes\left(\hat{\mathbb{G}}_{i}^{D}\oplus\mathcal{F}_{\Phi}(\hat{\mathbb{G}}_{i}^{D},e,t)\right)~{}~{}.$ (6)
3.3 Multi-Scale Bilateral Grid for Appearance Enhancement
Solving the photometric correction problem formulated in Eq. (4) fundamentally depends on the ability to effectively model the complex and spatially-varying photometric transformations present in real-world driving scenes. The function $\mathcal{E}(\cdot)$ in our formulation represents this crucial photometric enhancement process. To address this, we propose the Multi-Scale Bilateral Grid architecture, detailed in this section, as a novel grounded solution for approximating $\mathcal{E}(\cdot)$ and achieving high-quality, consistent appearance enhancement.
3.3.1 Multi-Scale Bilateral Grid Architecture
To address limitations of existing methods in handling diverse photometric variations in driving scenes, we propose a multi-scale bilateral grid framework. Our framework achieves this unification by employing a hierarchical pyramid of bilateral grids, organized across multiple scales. This multi-scale design is directly inspired by the nature of photometric inconsistencies in real-world environments, which range from global scene-level changes to highly localized variations (e.g., from overall lighting shifts to fine texture-level shadows). By utilizing this grid hierarchy, our framework aims to capture and effectively correct photometric variations at their corresponding scales, thus enabling a more comprehensive and spatially adaptable photometric correction.
The multi-scale bilateral grid transformation can be formulated as $I^{e}=\bar{A}\odot I^{r}$, where $\bar{A}$ represents a composite, scale-dependent photometric transformation. Crucially, this composite transformation is constructed hierarchically, combining transformations learned by individual bilateral grids at different scales, progressing from coarse to fine. This staged composition is key to achieving scale-dependent adaptation to photometric inconsistencies and allows us to move beyond the inherent limitations of single-scale methods. Further details are provided in the subsequent sections.
Our framework achieves this hierarchical representation using a three-level bilateral grid pyramid (Fig. 2): (1) Coarse Level ($2\times 2\times 1\times 12$ grid) captures global scene structure and approximate global appearance codes; (2) Intermediate Level ($4\times 4\times 2\times 12$ grid) represents regional features and capture mid-range photometric variations; (3) Fine Level ($8\times 8\times 4\times 12$ grid) encodes local details and approximate pixel-wise bilateral grids within the multi-scale framework.
At each level $l$, the grid tensor is defined as $\mathcal{A}^{(l)}\in\mathbb{R}^{H^{(l)}\times W^{(l)}\times D^{(l)}\times C^{(l)}}$, where Spatial Dimension $(H^{(l)},W^{(l)})$ represents spatial resolution (height, width) at level $l$, Guidance Dimension $D^{(l)}$ specifies the guidance intensity levels at level $l$ and Coefficient Channel $C=12$ represents a flattened $3\times 4$ affine color transformation matrix.
3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation chen2007real to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of chen2007real ; wang2024bilateral ; gharbi2017deep . This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid. For each level $l$ and pixel $(u,v)$, the luminance $d$ from $I^{g}$ is used to perform a slicing operation, querying the grid’s intensity dimension $D^{(l)}$ to locate an intensity bin. Affine transformation coefficients $\mathcal{A}^{(l)}_{i,j,k}$ around this bin are then combined via trilinear interpolation for a level-specific transformation $\bar{A}^{(l)}(u,v)$. The slicing operation retrieves per-pixel transformations from each grid level, written as:
$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{s}(i,j,k)~{}~{},$ (7)
To efficiently fuse transformations across scales, we utilize hierarchical function composition. A naive approach of full-resolution slicing at each scale is computationally expensive and redundant due to local photometric coherence. To address this, we employ downsampled guidance maps $I^{g{(l)}}$ for each scale $l$, performing slicing to obtain low-resolution coefficient fields, which are then upsampled. This approach explicitly links scale-aware guidance to patch operations-the reduced-resolution maps enable efficient spatial aggregation of photometric transformations while enhances efficiency (See ablation study in Tab. 5). The level-specific transformations $\mathcal{T}^{(l)}$, decomposed into linear matrices $\bar{M}^{(l)}$ and translations $\bar{T}^{(l)}$, are then sequentially composed from coarse to fine to produce the final composite transformation $\bar{A}$. The refined image $I^{e}(u,v)$ is obtained by applying $\bar{A}(u,v)$ to $I^{r}(u,v)$, with the fusion process expressed as:
$I^{e}=\mathcal{T}^{(L-1)}\circ\mathcal{T}^{(L-2)}\circ\cdots\circ\mathcal{T}^{(0)}(I^{r})~{}~{},$ (8)
This hierarchical fusion decomposes the photometric transformation into scale-dependent residual refinements. Coarse scales capture global transformations, intermediate scales refine regional variations, and fine scales address local details (Further elaboration in Sec. 4.3). Further details on the interpolation kernel and computational efficiency considerations are provided in Appendix A1.1.
Dataset Method Full Image Human Vehicle Geometry
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
Waymo sun2020scalability Baseline 28.92 0.833 25.24 0.681 23.50 0.697 1.482 2.785 0.540
Baseline w/AC 28.95 0.835 24.92 0.656 23.51 0.699 1.378 2.760 0.519
Baseline w/BG 28.19 0.831 24.79 0.664 23.09 0.693 1.523 2.798 0.492
Ours 29.23 0.836 25.55 0.683 23.50 0.707 0.989 2.744 0.477
NuScenes caesar2020nuscenes Baseline 26.37 0.837 26.17 0.774 26.49 0.827 1.458 3.420 0.110
Baseline w/AC 26.38 0.840 25.83 0.762 26.52 0.828 1.437 3.415 0.106
Baseline w/BG 25.98 0.837 26.00 0.771 26.10 0.820 1.380 3.390 0.097
Ours 27.69 0.847 26.53 0.777 27.31 0.834 1.161 3.340 0.059
Argoverse wilson2023argoverse Baseline 24.59 0.848 23.35 0.743 24.47 0.812 0.954 4.208 0.050
Baseline w/AC 24.58 0.848 23.22 0.738 24.44 0.811 0.959 4.215 0.051
Baseline w/BG 23.31 0.842 22.99 0.730 23.49 0.795 0.901 4.218 0.049
Ours 24.68 0.849 23.35 0.739 24.51 0.827 0.807 4.199 0.040
Pandaset xiao2021pandaset Baseline 30.20 0.903 29.79 0.830 23.27 0.776 0.503 2.874 0.018
Baseline w/AC 30.20 0.903 29.64 0.824 23.39 0.783 0.496 2.871 0.020
Baseline w/BG 29.73 0.904 29.90 0.812 22.22 0.763 0.484 2.867 0.013
Ours 30.75 0.906 30.18 0.830 23.49 0.784 0.453 2.852 0.011
Table 1: Comparison with SOTA for scene reconstruction. w/AC represents appearance codes while w/BG represents single bilateral grids.
4 Experiments
4.1 Experimental Setup
Datasets.  We evaluate on four autonomous driving datasets: Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations. Identical model hyperparameters are used across all datasets to test generalization capability. Technical specifications including camera counts and ego-vehicle view cropping details are provided in Appendix A2.
Baseline.  Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio chen2024omnire , ChatSim wei2024editable , StreetGS yan2024street and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids. This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
Method Reconstruction Novel View Synthesis Geometry
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
ChatSim wei2024editable 25.10 0.805 0.252 23.27 0.725 0.270 1.557 3.509 0.106
ChatSim (Ours) 27.04 0.819 0.231 24.67 0.735 0.249 1.236 3.412 0.053
StreetGS yan2024street 25.74 0.822 0.240 23.64 0.736 0.259 1.604 3.544 0.107
StreetGS (Ours) 27.90 0.836 0.219 25.10 0.747 0.238 1.272 3.458 0.055
Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the nuScenes caesar2020nuscenes scenarios.
Implementation details.  We evaluate our method on many autonomous driving datasets to demonstrate its robustness. For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis results between camera and LiDAR data. Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
Dataset Method PSNR $\uparrow$
Full Image Human Vehicle
Waymo sun2020scalability Baseline 26.40 23.91 21.25
Baseline w/AC 26.44 23.81 21.34
Baseline w/BG 21.51 22.91 19.82
Ours 26.55 23.97 21.36
NuScenes caesar2020nuscenes Baseline 23.74 23.55 22.80
Baseline w/AC 23.74 23.35 22.84
Baseline w/BG 23.60 23.63 22.79
Ours 24.64 23.75 23.22
Argoverse wilson2023argoverse Baseline 22.53 20.73 20.84
Baseline w/AC 22.51 20.70 20.80
Baseline w/BG 21.70 20.67 20.59
Ours 22.58 20.79 20.86
Pandaset xiao2021pandaset Baseline 27.49 26.94 21.02
Baseline w/AC 27.51 26.98 21.13
Baseline w/BG 27.38 27.08 20.77
Ours 27.89 27.22 21.32
Table 3: Comparison on Novel View Synthesis.
Training and Dynamic Rendering.
1) Training.  To optimize our multi-scale Gaussian scene representation, we employ a joint training strategy that minimizes a composite reconstruction loss :
$\mathcal{L}_{recon}=\lambda_{r}\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}~{}~{},$ (9)
Furthermore, we introduce two additional regularization terms to enhance image fidelity: Adaptive Total Variation Regularization. This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss. This loss applies an inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation. As visually demonstrated in Fig. 5, these two terms significantly improve image fidelity. Detailed explanations are provided in Appendix A1.
2) Rendering. For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions. Our proposed method first conducts a temporal proximity search to identify the most relevant grids. Following this, we perform scale-specific interpolation using the two nearest grids found. More comprehensive details regarding this interpolation approach are available in Appendix A1.
Figure 3: Visualizations of our method and baseline methods on Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset dataset.
Figure 4: Our proposed framework (c) outperforms appearance codes (a) and single bilateral grids (b) by simultaneously addressing optimization challenges and enhancing geometric modeling. This delivers significant geometric accuracy improvements (lower Chamfer Distance (CD) via multi-scale bilateral grid, reducing error and improving reconstruction) and reduced photometric inconsistency (fewer floaters) in dynamic, decoupled driving scenes. Yellow indicates a high lidar error, while Purple indicates a low lidar error.
4.2 Quantitative Evaluation
Reconstruction Quality.  We evaluate the reconstruction quality of our method using appearance and geometry metrics, with detailed results presented in Tab. 1.
(1) Appearance Fidelity. Our approach demonstrates superior performance in appearance-based metrics (PSNR and SSIM). As shown in Tab. 1, our method achieves the highest PSNR and SSIM scores for full image reconstruction across all tested datasets (Waymo, NuScenes, Argoverse, and Pandaset). Furthermore, when evaluating segmented dynamic regions such as humans and vehicles, our method largely surpasses the baselines, indicating its robustness in capturing both global scene appearance and fine-grained details of dynamic objects. For instance, on NuScenes, our PSNR for ’Vehicle’ is 27.31, compared to the best baseline score of 26.52. While there are isolated cases where a baseline performs comparably or slightly better (e.g., Human SSIM on Argoverse), our method generally offers the most consistent and highest quality appearance reconstruction.
(2) Geometric Accuracy. In terms of geometric fidelity, our method consistently outperforms all baselines across all datasets, as evidenced by the Chamfer Distance (CD), Root Mean Square Error (RMSE), and Depth L2 Error metrics in Tab. 1. For example, on the Waymo dataset, our method achieves a CD of 0.989, significantly lower than the best baseline score of 1.378. This superior geometric accuracy is partly attributed to our model’s improved handling of "floater" artifacts, leading to more precise LiDAR representations, as qualitatively illustrated in Fig. 4.
Novel View Synthesis.  The performance of our method in novel view synthesis is detailed in Tab. 3. Our approach consistently achieves the highest PSNR scores across all datasets for full images as well as for segmented human and vehicle categories. For example, on the NuScenes dataset, our method achieves a PSNR of 24.64 for the full image, 23.75 for humans, and 23.22 for vehicles, outperforming all baseline variations. This underscores the capability of our method to generate high-fidelity novel views from new camera poses.
Comparison with State-of-the-Art Methods. To further assess the benefits of our architectural contributions, we integrated our core components into two recent state-of-the-art methods: ChatSim wei2024editable and StreetGS yan2024street . The comparative results, averaged over NuScenes scenarios, are presented in Tab. 2.
When our enhancements are applied (denoted as "ChatSim (Ours)" and "StreetGS (Ours)"), both methods exhibit significant improvements across all evaluated aspects. For instance, integrating our approach with ChatSim improves its reconstruction PSNR from 25.10 to 27.04 and its novel view synthesis PSNR from 23.27 to 24.67. Similarly, StreetGS sees its reconstruction PSNR boosted from 25.74 to 27.90 and its Chamfer Distance reduced from 1.604 to 1.272. These results highlight the generalizability and effectiveness of our proposed techniques, demonstrating their capability to enhance the performance of existing strong baseline models for comprehensive scene understanding and representation.
4.3 Qualitative Evaluation
Rendering Results.  Fig. 3 presents visualization results of our method on scenes from Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
Figure 5: Visualizations of ablations on Circle Regularization loss and Adaptive TV loss.
Figure 6: Visualization of Affine Transformations and Distributions. The first three rows display the affine transformation results and corresponding distributions when applying the first level (scale 0), the first two levels (scales 0-1), and all three levels (scales 0-1-2) of our multi-scale approach. The final row compares these to results from a single bilateral grid, demonstrating that the multi-scale bilateral grid achieves a coarse-to-fine, smooth image enhancement with more diverse processing capabilities.
Affine Transformations Distributions.  According to Fig. 6, analyzing grayscale histograms of learned affine transformations reveals a key distinction between original single-scale and our multi-scale bilateral grids. For the original grid, each viewpoint’s histogram typically exhibits two distinct peaks. This bi-modal distribution suggests a focused, somewhat limited photometric adaptation. It implies that for each view, correction focuses on a few dominant transformations. More visualization results are shown at Fig. 3. A particularly interesting observation arises examining histograms of our multi-scale grid’s coarsest level. Intriguingly, these histograms also demonstrate peaky distributions, varying significantly across viewpoints. This isn’t contradictory, but insightful: even at its coarsest level, the grid learns view-dependent base appearances. The coarse grid captures dominant, scene-wide photometric transformations characteristic of each viewpoint. Peaks in coarse-level histograms likely represent these primary, global photometric adjustments, tailored per view.
Configuration Grid Params (M) Training Time (h) Testing time (FPS)
Baseline w/AC 0.072 1.93 53
Baseline w/BG 27.843 2.85 54 / 38†
Ours 3.969 2.10 54 / 42†
Table 4: Computational Efficiency Analysis. †  means using bilateral grids processing during rendering.
However, the key advantage is evident in the overlaid histogram, which aggregates distributions from all multi-scale levels and viewpoints. This histogram is notably flatter and more dispersed than those of the original grid, a characteristic directly resulting from hierarchical and residual refinement. Our framework first uses the coarse grid to establish a view-dependent base appearance. Subsequently, medium and fine grids learn residual transformations to correct coarse grid errors. Medium grids then refine regional photometric variations, while fine grids handle local, pixel-level details. This layered, residual approach significantly broadens the spectrum of learnable photometric transformations. The original grid’s two-peaked histograms signify its narrow adaptation. In contrast, the multi-scale version’s flatter histogram reveals its core strength: a diverse, generalizable, and robust representation of spatially-varying photometric transformations. This enhanced representation is crucial for consistent and high-quality renderings across diverse driving views.
4.4 Ablation Studies
As shown in Tab. 4, Tab. 5 and Fig. 5, we systematically evaluate component contributions through:
Settings Recon.$\uparrow$ NVS.$\uparrow$ CD $\downarrow$
   Full model 27.69 24.64 1.161
(a) Guidance Map Resolution
   (1,1,1) 27.62 24.34 1.271
   (2,2,1) 27.68 24.64 1.170
   (8,8,4) 27.52 24.53 1.187
   (16,16,8) 27.50 24.54 1.202
(b) Grid Size Combinations
   (4,4,2)+(8,8,4) 27.49 24.56 1.963
   (2,2,1)+(8,8,4) 27.57 24.57 1.940
   (2,2,1)+(4,4,2) 27.33 24.42 1.213
   (4,4,2)+(8,8,4)+(16,16,8) 27.31 24.46 2.297
(c) Loss Functions
   w/o circle regularization loss 27.47 24.52 1.164
   w/o adaptive total variance loss 27.64 24.61 1.169
Table 5: Ablations. Numbers in Guidance Map Resolution settings indicate downsample factors from coarse level to fine level, and numbers in Grid Size Combinations settings indicate grid size.
Guidance Map Resolution.   Varying guidance map resolution shows a trade-off in performance. Using the highest resolution (1,1,1) results in a slightly lower NVS PSNR compared to using (2,2,1) which achieves the highest NVS PSNR(24.64) in these ablations, demonstrating our patch-based approach balances efficiency and effectiveness. Grid Size Combinations.   Using different combinations of grid sizes significantly impacts geometry, as indicated by CD. Combining smaller grids like (2,2,1)+(4,4,2) results in a lower CD (1.213) than using combinations including larger grids or the combination with the largest grid (4,4,2)+(8,8,4)+(16,16,8). Loss Functions.   Ablating the circle regularization loss slightly reduces both Reconstruction PSNR (from 27.69 to 27.47) and NVS PSNR (from 24.64 to 24.52). Removing the adaptive total variance loss has a minimal impact on rendering result. Efficiency.   As shown in Tab. 4, while the ’Baseline w/BG’ dramatically increases parameters and training time, our method achieves a much more moderate increase in parameters and training time (to 2.10 hours) relative to the lightweight appearance codes. With multi-scale bilateral grid guided training, we found the rendering speed may improve due to the decreased number of unnecessary Gaussians.
5 Conclusion
We propose a multi-scale bilateral grid seamlessly integrating global appearance codes and pixel-wise bilateral grids to address photometric inconsistencies in autonomous driving neural rendering. Limitations. While our proposed multi-scale bilateral grid demonstrates significant improvements in handling photometric inconsistencies and enhancing geometric accuracy in driving scene reconstruction, certain limitations warrant discussion. (1) The computational remains higher than simpler appearance code techniques; this could be a factor in highly resource-constrained deployment scenarios. (2) while effective for many dynamic elements, modeling extremely complex, fast-moving, or highly non-rigid objects with complete fidelity remains an ongoing challenge.