### v4

# Unifying Appearance Codes and Bilateral Grids for Driving Scene Gaussian Splatting

Abstract
Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions. However, in real-world driving scenarios, it is challenging to guarantee perfect photometric consistency in acquired images. Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image. Recently, the bilateral grid was introduced to perform pixel-wise color mapping, but it is difficult to optimize and constrain effectively. In this paper, we propose a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids. We demonstrate that this approach significantly improves geometric accuracy in dynamic, decoupled autonomous driving scene reconstruction, outperforming both appearance codes and bilateral grids. This is crucial for autonomous driving, where accurate geometry is important for obstacle avoidance and control. Our method shows strong results across four datasets: Waymo, NuScenes, Argoverse, and PandaSet. We further demonstrate that the improvement in geometry is driven by the multi-scale bilateral grid, which effectively reduces floaters caused by photometric inconsistency. Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
Figure 1: Unifying appearance codes and bilateral grids. (a) Appearance codes rely on global affine transformations but have limited modeling capability. (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize. (c) The proposed multi-scale bilateral grid unifies both paradigms, enabling patch-wise transformations.
1 Introduction
Neural rendering techniques, such as NeRFs  Mildenhall et al. (2021); Wang et al. (2021); Yu et al. (2022); Liu et al. (2024); Yuan and Zhao (2024) and Gaussian Splatting (GS) Kerbl et al. (2023); Chen et al. (2024a); Cheng et al. (2024); Zhang et al. (2024b); Song et al. (2024); Ye et al. (2025); Shuai et al. (2025), have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency. However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably Martin-Brualla et al. (2021); Zhang et al. (2024a); Kulhanek et al. (2024); Wang et al. (2025); Gao et al. (2024). In autonomous driving applications Yan et al. (2024b); Wu et al. (2023); Tonderski et al. (2024); Chen et al. (2024b); Li et al. (2024b); Jin et al. (2024); Tian et al. (2023); Han et al. (2025); Yu et al. (2025), these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time. As illustrated in the left half of Teaser. 1, we observe that these variations—whether temporal (across different time steps) or inter-camera (across different viewpoints)—can introduce significant discrepancies in image appearance. These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
To address this issue, appearance codes Tonderski et al. (2024); Martin-Brualla et al. (2021); Chen et al. (2024b) have been introduced to encode per-image information and assist in correcting the photometric discrepancies. These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes. For example, in dynamic scenes, such as those encountered in autonomous driving environments, lighting and viewpoints can change across different camera angles and time frames. As shown in the right top of Teaser. 1, a global affine transformation based on a single appearance code may not be sufficient to handle such variations (In Teaser. 1 NuScenes (a), the fence is blurred).
Recent advancements Wang et al. (2024); Gharbi et al. (2017) have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments. However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser. 1 (b), bilateral grids fail to effectively optimize the reconstruction of large, complex scenes.). In this work, we propose a novel solution—a multi-scale bilateral grid—integrating the strengths of appearance codes and bilateral grids. As shown in Teaser. 1, this new approach allows for patch-wise transformations. Interestingly, in extreme cases, the multi-scale bilateral grid naturally converges to either the bilateral grid or appearance code, depending on the scale. At the finest scale, the multi-scale grid behaves like a traditional bilateral grid, performing pixel-wise transformations to adjust for local variations. In contrast, at the coarsest scale, it effectively reverts to a more global transformation, resembling the behavior of appearance codes. This flexible, scale-dependent approach offers the best of both worlds, providing localized fine-tuning where necessary, while maintaining global consistency when appropriate.
We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning. This is evidenced by our evaluation on multiple widely used datasets, including Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021). The corresponding Chamfer Distance (CD) values, displayed in the Teaser. 1, highlight a marked reduction in error and improvement in reconstruction quality. Additionally, our method reduces photometric inconsistency, mitigating the appearance of floaters and enhancing the overall realism of the scene.
The contributions of this paper are as follows: First, we introduce a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids, transitioning to either of the two paradigms in extreme cases, thus enhancing both modeling capability and optimization efficiency. Second, we show that by addressing photometric inconsistencies, it improves the geometric accuracy of dynamic, decoupled driving scene reconstructions. Third, we provide extensive benchmarking across four widely used datasets—Waymo, NuScenes, Argoverse, and PandaSet—where our method outperforms previous approaches, showcasing notable improvements in both qualitative and quantitative results.
Figure 2: Overview of our method. We unify appearance codes with multi-scale bilateral grids. A coarse rendering from the Gaussian scene graph is refined by multi-scale bilateral grids to model per-pixel color with a luminance-guided slice-and-fuse pipeline.
2 Related Works
2.1 Bilateral Grids and Appearance Codes
Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving. The bilateral grids, first introduced for real-time edge-aware image processing Chen et al. (2007), enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques Gharbi et al. (2017); Märki et al. (2016); Chen et al. (2016); Xu et al. (2021); Zheng et al. (2021). Extensions like deep bilateral learning Gharbi et al. (2017) further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices. Similarly, bilateral-guided upsampling Chen et al. (2016) and cost volume refinement Xu et al. (2021) have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations. In the context of neural rendering, methods such as NeRF Tancik et al. (2022); Jang and Agapito (2021); Martin-Brualla et al. (2021) and Gaussian Splatting (GS) Zhang et al. (2024a); Kulhanek et al. (2024) have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders Martin-Brualla et al. (2021); Zhang et al. (2024a) remains a challenge. Techniques like bilateral guided radiance field processing Wang et al. (2024) address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space. WildGaussians Kulhanek et al. (2024) integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes. Cross-Ray NeRF Yang et al. (2023b) employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections. Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability. For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations Durand and Dorsey (2002). Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively Paris and Durand (2006). To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
2.2 Autonomous Driving Simulation
Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios Wu et al. (2023); Okae et al. (2021); Yan et al. (2024a); Lindström et al. (2024); Tao et al. (2024b); Li et al. (2024b); Yang et al. (2024a); Gao et al. (2023). Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation Wu et al. (2023); Li et al. (2024a); Zhou et al. (2024b); Tonderski et al. (2024); Yan et al. (2024a); Yang et al. (2023a, 2025b); Xiao et al. (2025). For instance, MARS Wu et al. (2023) employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian Yan et al. (2024a) achieves real-time urban scene rendering using explicit 3D Gaussian representations. DrivingGaussian Zhou et al. (2024b) enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency. NeuRAD Tonderski et al. (2024) integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis. Holistic scene understanding and editing have also seen progress. HUGS Zhou et al. (2024a) combines static and dynamic 3D Gaussians for real-time semantic parsing. ChatSim Wei et al. (2024) enables language-driven scene editing with external asset integration. LiDAR data integration has advanced through methods like LiDAR-NeRF Tao et al. (2024a), which uses structural regularization for low-texture regions. HO-Gaussian Li et al. (2024c) merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion dependencies.
Despite progress, gaps persist between simulated and real-world data Lindström et al. (2024); Tao et al. (2024b); Yang et al. (2024b, 2025a). AlignMiF Tao et al. (2024b) addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS Nguyen et al. (2024) decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts. Innovations in large-scale generation include InfiniCube Lu et al. (2024), which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire Chen et al. (2024b), which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations. Furthermore, recent works like GaussianPro Cheng et al. (2024) and SplatAD Hess et al. (2024) explore to refine 3D GS for real-time rendering of dynamic scenes. GaussianPro Cheng et al. (2024) introduces progressive propagation for texture-less surfaces, and SplatAD Hess et al. (2024) models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity). In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
3 Methodology
3.1 Problem Formulation
We aim to reconstruct a 3D scene representation $\mathbb{G}$ from multi-view images and LiDAR depth maps, commonly available in autonomous driving datasets. Given a set of images $\{I_{c,t}\}$ and corresponding depths $\{D_{c,t}\}$ captured by $N$ cameras over $T$ time steps, we formulate our objective as minimizing the discrepancy between rendered and observed images and depths:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|I_{c,t}^{r}-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~~,$ (1)
where $I_{c,t}^{r}$ and $D_{c,t}^{r}$ are the rendered appearance and geometry of the scene $\mathbb{G}$ from camera $c$ at time $t$, and $\lambda_{D}$ is a weight balancing the depth term.
However, dynamic driving scenes and varying camera properties cause inconsistent appearances, leading to geometric and texture artifacts. To address inconsistency, we decompose each image $I_{c,t}$ into consistent $\mathcal{C}_{c,t}$ (e.g., intrinsic scene colors and constant sensor adjustments like normalization exposure), and non-consistent $\mathcal{N}_{c,t}$ (e.g., components-varying lighting conditions, different camera settings and diverse image ISP effects):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{N}_{c,t}~~,$ (2)
We model $\mathcal{N}_{c,t}$ as a non-linear transformation $\mathcal{F}(\mathcal{C}_{c,t})$. Substituting into Eq. (2):
$I_{c,t}=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})~~,$ (3)
Let $\mathcal{E}(\mathcal{C}_{c,t})=\mathcal{C}_{c,t}+\mathcal{F}(\mathcal{C}_{c,t})$, we can reformulate the optimization objective in Eq. (1) as:
$\min_{\mathbb{G}}\sum_{c,t}(\left\|\mathcal{E}(\mathcal{C}_{c,t}^{r})-I_{c,t}\right\|_{2}^{2}+\lambda_{D}\left\|D_{c,t}^{r}-D_{c,t}\right\|_{2}^{2})~~,$ (4)
This reformulation links photometric and geometric consistency optimization within a joint objective, guiding $\mathcal{C}_{c,t}^{r}$ adjustments by geometric constraints ($D_{c,t}^{r}$). This mitigates texture-geometry ambiguities (e.g., shadows on road surfaces misinterpretable as geometry changes). $\mathcal{F}(\mathcal{C}_{c,t})$ models transient appearance, while $\mathcal{C}_{c,t}^{r}$ enforces consistent scene properties. Experiments (Sec. 4, Tab. 1) show this joint optimization reduces both photometric error and geometric drift in comparison to baselines.
3.2 Gaussian Splatting for Autonomous Driving Environments
Following prior methodologies Chen et al. (2024b); Tonderski et al. (2024), we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models. Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations. We derive the transformation matrixes from existing object detection pipelines or ground-truth annotations.
For the sky, we use an environment map to model sky color based on viewing direction, while static background is represented as a set of semi-transparent 3D Gaussians. Each Gaussian is characterized by a learnable opacity parameter $o\in(0,1)$, a mean position $\mu\in\mathbb{R}^{3}$, and an anisotropic covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ which is parameterized by a scale vector $S\in\mathbb{R}^{3}$ and a rotation quaternion $q\in\mathbb{R}^{4}$. Additionally, Spherical harmonics coefficients $c\in\mathbb{R}^{F}$ are used to model appearance.
For the moving object models, we further distinguish between non-deformable (e.g., vehicles) and deformable objects (e.g., pedestrians). Non-deformable objects $\{\mathbb{G}_{i}^{N}|i\in\{1,\dots,n_{n}\}\}$ are optimized in their local coordinate space and transformed into the global world space via their pose $T$:
$\mathbb{G}_{i}^{N}=T\otimes\hat{\mathbb{G}}_{i}^{N}~~,$ (5)
For deformable objects $\{\mathbb{G}_{i}^{D}|i\in\{1,\dots,n_{d}\}\}$, we employ a deformation network $\mathcal{F}_{\Phi}$ (parameterized by $\Phi$) to adapt the Gaussian representation based on latent variable $e$ and time $t$:
$\mathbb{G}_{i}^{D}=T\otimes\left(\hat{\mathbb{G}}_{i}^{D}\oplus\mathcal{F}_{\Phi}(\hat{\mathbb{G}}_{i}^{D},e,t)\right)~~.$ (6)
3.3 Multi-Scale Bilateral Grid for Appearance Enhancement
Solving the photometric correction problem formulated in Eq. (4) fundamentally depends on the ability to effectively model the complex and spatially-varying photometric transformations present in real-world driving scenes. The function $\mathcal{E}(\cdot)$ in our formulation represents this crucial photometric enhancement process. To address this, we propose the Multi-Scale Bilateral Grid architecture, detailed in this section, as a novel grounded solution for approximating $\mathcal{E}(\cdot)$ and achieving high-quality, consistent appearance enhancement.
3.3.1 Multi-Scale Bilateral Grid Architecture
To address limitations of existing methods in handling diverse photometric variations in driving scenes, we propose a multi-scale bilateral grid framework. Our framework achieves this unification by employing a hierarchical pyramid of bilateral grids, organized across multiple scales. This multi-scale design is directly inspired by the nature of photometric inconsistencies in real-world environments, which range from global scene-level changes to highly localized variations (e.g., from overall lighting shifts to fine texture-level shadows). By utilizing this grid hierarchy, our framework aims to capture and effectively correct photometric variations at their corresponding scales, thus enabling a more comprehensive and spatially adaptable photometric correction.
The multi-scale bilateral grid transformation can be formulated as $I^{e}=\bar{A}\odot I^{r}$, where $\bar{A}$ represents a composite, scale-dependent photometric transformation. Crucially, this composite transformation is constructed hierarchically, combining transformations learned by individual bilateral grids at different scales, progressing from coarse to fine. This staged composition is key to achieving scale-dependent adaptation to photometric inconsistencies and allows us to move beyond the inherent limitations of single-scale methods. Further details are provided in the subsequent sections.
Our framework achieves this hierarchical representation using a three-level bilateral grid pyramid (Fig. 2): (1) Coarse Level ($2\times 2\times 1\times 12$ grid) captures global scene structure and approximate global appearance codes; (2) Intermediate Level ($4\times 4\times 2\times 12$ grid) represents regional features and capture mid-range photometric variations; (3) Fine Level ($8\times 8\times 4\times 12$ grid) encodes local details and approximate pixel-wise bilateral grids within the multi-scale framework.
At each level $l$, the grid tensor is defined as $\mathcal{A}^{(l)}\in\mathbb{R}^{H^{(l)}\times W^{(l)}\times D^{(l)}\times C^{(l)}}$, where Spatial Dimension $(H^{(l)},W^{(l)})$ represents spatial resolution (height, width) at level $l$, Guidance Dimension $D^{(l)}$ specifies the guidance intensity levels at level $l$ and Coefficient Channel $C=12$ represents a flattened $3\times 4$ affine color transformation matrix.
3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation Chen et al. (2007) to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of Chen et al. (2007); Wang et al. (2024); Gharbi et al. (2017). This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid. For each level $l$ and pixel $(u,v)$, the luminance $d$ from $I^{g}$ is used to perform a slicing operation, querying the grid’s intensity dimension $D^{(l)}$ to locate an intensity bin. Affine transformation coefficients $\mathcal{A}^{(l)}_{i,j,k}$ around this bin are then combined via trilinear interpolation for a level-specific transformation $\bar{A}^{(l)}(u,v)$. The slicing operation retrieves per-pixel transformations from each grid level, written as:
$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{(l)}(i,j,k)~~,$ (7)
To efficiently fuse transformations across scales, we utilize hierarchical function composition. A naive approach of full-resolution slicing at each scale is computationally expensive and redundant due to local photometric coherence. To address this, we employ downsampled guidance maps $I^{g{(l)}}$ for each scale $l$, performing slicing to obtain low-resolution coefficient fields, which are then upsampled. This approach explicitly links scale-aware guidance to patch operations-the reduced-resolution maps enable efficient spatial aggregation of photometric transformations while enhances efficiency (See ablation study in Tab. 5). The level-specific transformations $\mathcal{T}^{(l)}$, decomposed into linear matrices $\bar{M}^{(l)}$ and translations $\bar{T}^{(l)}$, are then sequentially composed from coarse to fine to produce the final composite transformation $\bar{A}$. The refined image $I^{e}(u,v)$ is obtained by applying $\bar{A}(u,v)$ to $I^{r}(u,v)$, with the fusion process expressed as:
$I^{e}=\mathcal{T}^{(L-1)}\circ\mathcal{T}^{(L-2)}\circ\cdots\circ\mathcal{T}^{(0)}(I^{r})~~,$ (8)
This hierarchical fusion decomposes the photometric transformation into scale-dependent residual refinements. Coarse scales capture global transformations, intermediate scales refine regional variations, and fine scales address local details (Further elaboration in Sec. 4.3). Further details on the interpolation kernel and computational efficiency considerations are provided in Appendix A1.1.
Table 1: Ablation studies and comparisons on four large-scale driving datasets. Our proposed method consistently outperforms OmniRe across all metrics and datasets. w/AC denotes appearance codes; w/BG denotes single bilateral grids (size of $16\times 16\times 8$).
Dataset Method Reconstruction Novel View Synthesis Geometry
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
Waymo OmniRe $28.92$ $0.833$ 0.295 $26.40$ 0.761 0.311 $1.482$ $2.785$ $0.540$
OmniRe w/AC $28.95$ $0.835$ 0.293 $26.44$ 0.759 0.315 $1.378$ $2.760$ $0.519$
OmniRe w/BG $28.19$ $0.831$ 0.292 $21.51$ 0.743 0.333 $1.523$ $2.798$ $0.492$
Ours 29.23 0.836 0.289 26.55 0.762 0.310 0.989 2.744 0.477
NuScenes OmniRe $26.37$ $0.837$ 0.209 $23.74$ 0.733 0.232 $1.458$ $3.420$ $0.110$
OmniRe w/AC $26.38$ $0.840$ 0.204 $23.74$ 0.732 0.229 $1.437$ $3.415$ $0.106$
OmniRe w/BG $25.98$ $0.837$ 0.209 $23.60$ 0.705 0.262 $1.380$ $3.390$ $0.097$
Ours 27.69 0.847 0.193 24.64 0.739 0.216 1.161 3.340 0.059
Argoverse OmniRe $24.59$ $0.848$ 0.202 $22.53$ 0.755 0.220 $0.954$ $4.208$ $0.050$
OmniRe w/AC $24.58$ $0.848$ 0.201 $22.51$ 0.756 0.219 $0.959$ $4.215$ $0.051$
OmniRe w/BG $23.31$ $0.842$ 0.216 $21.70$ 0.725 0.254 $0.901$ $4.218$ $0.049$
Ours 24.68 0.849 0.200 22.58 0.756 0.217 0.807 4.199 0.040
PandaSet OmniRe $30.20$ $0.903$ 0.219 $27.49$ 0.835 0.240 $0.503$ $2.874$ $0.018$
OmniRe w/AC $30.20$ $0.903$ 0.220 $27.51$ 0.841 0.242 $0.496$ $2.871$ $0.020$
OmniRe w/BG $29.73$ $0.904$ 0.220 $27.38$ 0.830 0.246 $0.484$ $2.867$ $0.013$
Ours 30.75 0.906 0.213 27.89 0.847 0.235 0.453 2.852 0.011
4 Experiments
4.1 Experimental Setup
Datasets.  We evaluate on four autonomous driving datasets: Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021), selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations. Identical model hyperparameters are used across all datasets to test generalization capability. Technical specifications including camera counts and ego-vehicle view cropping details are provided in Appendix A2.
Baseline.  Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio Chen et al. (2024b), ChatSim Wei et al. (2024), StreetGS Yan et al. (2024a) and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids. This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the NuScenes Caesar et al. (2020) scenarios.
Method Reconstruction Novel View Synthesis Geometry
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
ChatSim Wei et al. (2024) $25.10$ $0.805$ $0.252$ $23.27$ $0.725$ $0.270$ $1.557$ $3.509$ $0.106$
ChatSim (Ours) 27.04 0.819 0.231 24.67 0.735 0.249 1.236 3.412 0.053
StreetGS Yan et al. (2024a) $25.74$ $0.822$ $0.240$ $23.64$ $0.736$ $0.259$ $1.604$ $3.544$ $0.107$
StreetGS (Ours) 27.90 0.836 0.219 25.10 0.747 0.238 1.272 3.458 0.055
Implementation details.  We evaluate our method on many autonomous driving datasets to demonstrate its robustness. For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis (NVS) results between camera and LiDAR data. Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
Training and Dynamic Rendering.
1) Training.  To optimize our multi-scale Gaussian scene representation, we employ a joint training strategy that minimizes a composite reconstruction loss :
$\mathcal{L}_{recon}=\lambda_{r}\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}~~,$ (9)
Furthermore, we introduce two regularization terms to enhance image fidelity: Adaptive Total Variation Regularization. This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss. This loss applies inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation. As visually demonstrated in Fig. 5, these terms effectively improve image fidelity. Detailed are provided in Appendix A1.
2) Rendering. For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions. We first conducts a temporal proximity search to identify the most relevant grids. Following this, we perform scale-specific interpolation using the two nearest grids found.
Table 3: Evaluation on Challenging Scene Subsets. Results on systematically selected extreme cases. Our method shows substantially improvements when photometric inconsistencies are severe.
Scene Type Method Reconstruction Novel View Synthesis Geometry
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
Challenging Scenes (18 scenes) OmniRe $26.84$ $0.847$ 0.239 $24.46$ $0.758$ 0.263 $0.650$ $3.176$ $0.063$
Ours 28.13 0.856 0.223 25.17 0.763 0.250 0.523 3.120 0.023
Night Scenes (4 scenes) OmniRe $28.34$ $0.805$ 0.338 $26.54$ $0.753$ 0.360 $0.586$ $2.767$ $0.041$
Ours 29.13 0.812 0.330 27.21 0.762 0.349 0.500 2.731 0.016
Sun Flare Scenes (5 scenes) OmniRe $26.68$ $0.866$ 0.195 $24.98$ $0.796$ 0.212 $0.781$ $3.073$ $0.079$
Ours 28.65 0.879 0.173 25.84 0.803 0.192 0.593 2.994 0.022
Figure 3: Qualitative comparison across datasets. Our method versus baselines on Waymo, NuScenes, Argoverse, and PandaSet.
Figure 4: Photometric consistency improves geometry. Our framework (c) outperforms appearance codes (a) and single bilateral grids (b) by addressing optimization challenges and enhancing geometric modeling. This yields lower Chamfer Distance and fewer floaters in dynamic, decoupled driving scenes. Yellow indicates high LiDAR error, while Purple indicates low LiDAR error.
4.2 Quantitative Evaluation
Reconstruction and Geometric Quality. We present a comprehensive quantitative analysis in Tab. 1, comparing our full model against the OmniRe and its variants with appearance codes (w/AC) and a single bilateral grid (w/BG). Our method consistently sets a new state-of-the-art across all four datasets in both appearance and, most critically, geometric metrics.
In terms of appearance fidelity, our model achieves the highest PSNR and SSIM scores. For example, on NuScenes, our method reaches a PSNR of 27.69, a significant jump from 26.37 (OmniRe) and 26.38 (OmniRe w/AC). This demonstrates the superior ability of the multi-scale grid to model complex photometric variations.
However, the most striking result is the substantial improvement in geometric accuracy. Our method drastically reduces the Chamfer Distance, a key indicator of geometric fidelity. On the Waymo dataset, our model hits a CD of 0.989, outperforming the best baseline (OmniRe w/AC, 1.378) by a remarkable 28.2%. This trend holds across all datasets, with significant reductions in Chamfer Distance (CD), Root Mean Squared Error (RMSE), and Median Squared Depth Error (Depth). This result provides strong evidence for our central hypothesis: by accurately resolving photometric ambiguities, our multi-scale approach effectively eliminates geometric artifacts like "floaters" and produces a much cleaner, more accurate 3D reconstruction. The single bilateral grid (w/BG), despite its high expressiveness, often fails to converge and even degrades performance, highlighting the optimization challenges that our multi-scale, coarse-to-fine framework successfully overcomes.
NVS. As shown in Tab. 1, our model’s superior reconstruction quality translates directly to improved NVS. Our method consistently outperforms all baselines in NVS metrics across the four datasets. For instance, on PandaSet, we achieve a PSNR of 27.89, surpassing the next best (OmniRe w/AC at 27.51). This indicates that the cleaner geometry and more consistent appearance learned by our model generalize better to unseen viewpoints, producing higher-fidelity renderings.
Generalization to Other Methods. To demonstrate the broad applicability of our approach, we integrated our multi-scale bilateral grid into two other state-of-the-art methods, ChatSim Wei et al. (2024) and StreetGS Yan et al. (2024a). The results on NuScenes, presented in Tab. 2, are compelling. Our module brings substantial gains to both frameworks. For StreetGS, incorporating our method boosts the reconstruction PSNR from 25.74 to 27.90 and slashes the Chamfer Distance from 1.604 to 1.272. These significant improvements underscore that our multi-scale grid is not just a bespoke enhancement for one architecture but a generalizable and powerful module for improving photometric and geometric consistency in Gaussian Splatting-based scene reconstruction.
Performance on Challenging Scenarios. To validate robustness, we curated a subset of 18 challenging scenarios with extreme photometric inconsistencies. As shown in Table 3, our method’s advantages are substantially more pronounced here than on the standard datasets. Across this entire subset, our model achieves a +1.29 dB PSNR gain. This improvement is even more significant in specific difficult sub-categories; for instance, in "Sun Flare Scenes," the gain reaches +1.97 dB. This demonstrates our model’s strong effectiveness in resolving severe, real-world edge cases.
4.3 Qualitative Evaluation
Rendering Results.  Fig. 3 and Fig. 4 present visualization results of our method on scenes from Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021), demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
Figure 5: Ablation visualizations. Effects of circle regularization and adaptive total variation losses.
Figure 6: Affine transformation visualization. The first three rows show transformations and distributions for the first level, first two levels, and all three levels of our multi-scale approach. The final row compares against a single bilateral grid, illustrating a coarse-to-fine enhancement with richer transformation diversity.
Configuration Grid Params (M) Training Time (h) Testing Time (FPS)
OmniRe w/AC 0.072 1.93 53
OmniRe w/BG 27.843 2.85 54 / 38†
Ours 3.969 2.10 54 / 42†
Table 4: Computational Efficiency Analysis. † denotes rendering with bilateral grid processing active.
Analysis of Learned Affine Transformations. To gain deeper insight into how our model functions, we analyze the distribution of the learned affine transformations in Fig. 6. The histograms for the original single-scale bilateral grid (BG) exhibit a distinct bi-modal, peaky distribution for each viewpoint. This suggests that the single grid learns a limited set of dominant photometric corrections for each view, struggling to capture the full spectrum of variations. In stark contrast, the aggregated histogram from our multi-scale grid (across all levels and viewpoints) is significantly flatter and more dispersed. This is a direct visualization of our framework’s core strength. The coarse-level grid first establishes a view-dependent baseline appearance, capturing the dominant global photometric shift (interestingly, its own histogram is also peaky, but varies across views). Then, the medium and fine grids learn residual transformations in a coarse-to-fine manner, correcting the errors from the preceding levels. This hierarchical, residual refinement process allows the model to represent a much broader and more diverse range of photometric transformations. The flatter histogram is empirical evidence of this enhanced representational power, which is crucial for achieving consistent, high-quality renderings across the diverse and challenging views encountered in autonomous driving.
4.4 Ablation Studies
As shown in Tab. 4, Tab. 5 and Fig. 5, we systematically evaluate component contributions through:
[…]
  Full model 27.69 24.64 1.161
(a) Guidance Map Resolution
  (1,1,1) 27.62 24.34 1.271
  (2,2,1) 27.68 24.64 1.170
  (8,8,4) 27.52 24.53 1.187
  (16,16,8) 27.50 24.54 1.202
(b) Grid Size Combinations
  Single Grid (8,8,4) 26.56 23.90 1.376
  (4,4,2)+(8,8,4) 27.49 24.56 1.963
  (2,2,1)+(8,8,4) 27.57 24.57 1.940
  (2,2,1)+(4,4,2) 27.33 24.42 1.213
(c) Loss Functions
  w/o circle regularization loss 27.47 24.52 1.164
  w/o adaptive total variance loss 27.64 24.61 1.169
Table 5: Ablations. Numbers in Guidance Map Resolution indicate downsample factors from coarse to fine level, and numbers in Grid Size Combinations indicate grid size.
Guidance Map Resolution. We test different downsampling factors for the guidance map at each grid level. The results in Tab. 5(a) show that a moderate downsampling of (2,2,1) from coarse to fine yields the best balance, achieving the highest NVS PSNR (24.64). Using the full resolution (1,1,1) slightly degrades performance, likely due to overfitting to noise, validating our patch-based aggregation strategy.
Grid Size Combinations. The combination of grid sizes is crucial. As shown in Tab. 5(b), removing the coarsest grid level (e.g., using only (4,4,2)+(8,8,4)) leads to a dramatic increase in CD, highlighting the importance of the coarse grid for establishing global geometric consistency. Crucially, a baseline with only a single fine-resolution grid (8,8,4) performs poorly, confirming that finer grids fail to converge without the stable, coarse-level initialization our method provides.
Loss Functions. Ablating our proposed regularization losses in Tab. 5(c) shows their contribution. Removing the circle regularization loss slightly degrades both reconstruction and NVS performance. The effect of the adaptive total variance loss is more subtle in metrics but visibly improves rendering quality by reducing noise, as shown in Fig. 5.
Efficiency. As detailed in Tab. 4, our method strikes an effective balance between performance and computational cost. The single large bilateral grid (w/BG) leads to a massive parameter count (27.8M) and the longest training time (2.85h). In contrast, our multi-scale approach requires only 3.9M parameters and a modest 9% increase in training time (2.10h vs. 1.93h) compared to the appearance code baseline, while delivering substantially better results. During inference, our method maintains high frame rates, demonstrating its practicality.
5 Conclusion
We introduced a novel multi-scale bilateral grid that unifies global appearance codes and pixel-wise bilateral grids into a single, hierarchical framework. Our extensive experiments demonstrate that this approach not only effectively models and corrects complex photometric inconsistencies in autonomous driving scenes but, more importantly, leads to a significant and crucial improvement in the geometric accuracy of the final reconstruction. By establishing this strong link between advanced appearance modeling and geometric fidelity, our work paves the way for more robust and reliable neural rendering systems in safety-critical, real-world applications.
Limitations. While our method shows significant improvements, some limitations remain. (1) The computational overhead, though much lower than a single large bilateral grid, is still higher than simple appearance codes, which could be a consideration for resource-constrained scenarios. (2) Modeling extremely fast-moving or highly non-rigid objects where LiDAR and camera data misalign remains an open challenge.
Appendix A Additional Implementation Details
A.1 Loss Functions and Optimization
We jointly optimize our multi-scale Gaussian scene representation by minimizing a composite loss function $\mathcal{L}_{\text{total}}$:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{recon}}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}+\lambda_{\text{circle}}\mathcal{L}_{\text{circle}}$ (10)
where $\mathcal{L}_{\text{recon}}$ is the primary reconstruction loss, and $\mathcal{L}_{\text{TV}}$ and $\mathcal{L}_{\text{circle}}$ are regularization terms.
A.1.1 Reconstruction Loss
The core reconstruction loss, $\mathcal{L}_{\text{recon}}$, drives the accurate reproduction of both RGB and depth information:
$\mathcal{L}_{\text{recon}}=\lambda_{r}\mathcal{L}_{1}+(1-\lambda_{r})\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}$ (11)
Here, $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ measure the image-space difference. $\mathcal{L}_{d}$ represents the loss between the rendered depth and the ground truth LiDAR depth. $\mathcal{L}_{o}$ is an opacity regularization term that encourages alignment with a non-sky mask.
A.1.2 Adaptive Total Variation (TV) Regularization
To encourage smoothness and reduce noise in the optimized grid representations, we incorporate a TV regularization term, $\mathcal{L}_{\text{TV}}$, at each level $l$ of the multi-scale representation:
$\mathcal{L}_{\text{TV}}=\sum_{l}k^{(l)}\cdot\frac{1}{|\mathcal{A}^{(l)}|}\sum_{i,j,k}\sum_{\mathbb{D}\in\{x,y,z\}}\left\|\Delta_{\mathbb{D}}\mathcal{A}^{(l)}(i,j,k)\right\|_{2}^{2}$ (12)
The adaptive weight $k^{(l)}$ for each level $l$ is proportional to the grid size, applying stronger smoothing to finer, higher-resolution grids and lighter regularization to coarser grids.
$k^{(l)}=a\sqrt{H^{(l)}\cdot W^{(l)}\cdot D^{(l)}}+b$ (13)
A.1.3 Circle Regularization for Photometric Consistency
To constrain the noise introduced by photometric corrections and prevent overly aggressive alterations, we introduce a circle regularization loss, $\mathcal{L}_{\text{circle}}$:
$\mathcal{L}_{\text{circle}}=\sum_{(u,v)\in\mathcal{S}}\left\|I^{r}(u,v)-\bar{A}^{-1}(I^{gt}(u,v))\right\|_{2}^{2}$ (14)
where $\mathcal{S}$ denotes the set of pixel coordinates. This loss encourages the rendered image $I^{r}$ to be reconstructible from the ground truth image $I^{gt}$ via an inverse appearance transformation $\bar{A}^{-1}$.
A.1.4 Coarse-to-Fine Optimization Strategy
We employ a coarse-to-fine optimization strategy by utilizing level-dependent learning rates (e.g., $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$ from coarse to fine). Coarser grids, assigned higher learning rates, rapidly learn the global scene illumination, while finer grids, with lower learning rates, hierarchically refine high-frequency photometric details. This staged optimization enhances stability.
A.2 Dynamic Rendering for Real-World ISP Adaptation
To bridge the domain gap between the fixed ISP used during training and dynamic real-world ISP pipelines encountered during inference, we employ an interpolation strategy. For a novel test image with timestamp $t_{\text{novel}}$, we proceed as follows:
  1. 1.
    Temporal Proximity Search: We identify the two training timestamps from the same camera, $t_{1}$ and $t_{2}$, that are temporally closest to $t_{\text{novel}}$ (such that $t_{1}\leq t_{\text{novel}}\leq t_{2}$).
  2. 2.
    Scale-Specific Grid Interpolation: We perform linear interpolation specifically on the coarse and medium-scale bilateral grids ($\mathcal{A}^{(0)}$, $\mathcal{A}^{(1)}$) derived from these two timestamps.
The interpolated grid $\hat{\mathcal{A}}^{(l)}$ for $l\in\{0,1\}$ is formulated as:
$\hat{\mathcal{A}}^{(l)}=\omega\mathcal{A}^{(l)}_{t_{1}}+(1-\omega)\mathcal{A}^{(l)}_{t_{2}}$ (15)
The temporal interpolation weight, $\omega$, is determined by proximity:
$\omega=\frac{t_{2}-t_{novel}}{t_{2}-t_{1}}$ (16)
The fine-scale grid ($l=2$) is not used during interpolation, as it is primarily intended to capture scene-intrinsic details rather than global ISP variations.
Appendix B Dataset Details
This section provides dataset-specific details regarding our evaluation protocol, including sequence IDs for reproducibility.
B.1 Waymo Open Dataset Sun et al. (2020)
We use all five cameras and all LiDAR sensors. We conduct our experiments on the following 6 sequences, selected according to Tonderski et al. (2024); Wei et al. (2024):
  • segment-10017090168044687777
  • segment-10061305430875486848
  • segment-10584247114982259878
  • segment-15090871771939393635
  • segment-4458730539804900192
  • segment-5835049423600303130
B.2 NuScenes Caesar et al. (2020)
We utilize all six available cameras and all LiDAR sensors. We select the following 8 sequences: 152, 164, 171, 200, 209, 359, 529, 916, which is an extension of the dataset used by Tonderski et al. (2024). To address ego-vehicle visibility, we crop the bottom 80 pixels from the back camera images.
B.3 PandaSet Xiao et al. (2021)
We utilize six cameras and one LiDAR unit. We specifically evaluate on the following 10 challenging nighttime sequences: 063, 066, 070, 073, 074, 077, 078, 079, 088, 149. To mitigate ego-vehicle artifacts, we apply a bottom crop of 260 pixels to the back camera images.
B.4 Argoverse2 Wilson et al. (2023)
We leverage the seven ring cameras and both LiDAR sensors available in the dataset. We conduct our experiments on the following 9 sequences, in line with Tonderski et al. (2024):
  • 05fa5048-f355-3274-b565-c0ddc547b315
  • 0b86f508-5df9-4a46-bc59-5b9536dbde9f
  • 185d3943-dd15-397a-8b2e-69cd86628fb7
  • 25e5c600-36fe-3245-9cc0-40ef91620c22
  • 27be7d34-ecb4-377b-8477-ccfd7cf4d0bc
  • 280269f9-6111-311d-b351-ce9f63f88c81
  • 2f2321d2-7912-3567-a789-25e46a145bda
  • 44adf4c4-6064-362f-94d3-323ed42cfda9
  • 5589de60-1727-3e3f-9423-33437fc5da4b
To minimize ego-vehicle interference, we apply a bottom crop of 250 pixels to the front center, rear left, and rear right camera views.
Appendix C Additional Quantitative Results
This section presents a comprehensive breakdown of our model’s performance, including per-scene metrics, integration results with SOTA methods, and robustness checks on challenging subsets.
C.1 Per-Scene Geometry and Appearance Evaluation
We provide detailed metrics for geometry (Tab. 6 - 9) and appearance (Tab. 10 - 13).
C.2 Integration Evaluation with SOTA Methods
Tab. 14 and 15 provide comparisons with SOTA methods like ChatSim Wei et al. (2024) and StreetGS Yan et al. (2024a) on the NuScenes dataset.
C.3 Performance on Challenging Scenarios
To validate robustness, we curated subsets of challenging scenarios with extreme photometric inconsistencies (e.g., night scenes, sun glare, rain, and complex reflections). As shown in Tables 16, 17, and 18, our method’s performance gap widens significantly compared to all baselines (OmniRe, w/ AC, w/ BG) in these difficult conditions, demonstrating the superior robustness of our multi-scale design.
  • Challenging Argoverse (6 scenes): 49e970c4 (Extreme Lighting, Overexposure, Lens Flare), 8184872e (Complex Shadows, Direct Sunlight), 91923e20 (Extreme Lighting, Camera Artifacts), a8a2fbc2 (Low-Angle Sun), b403f8a3 (Overpass Structure, Extreme Lighting Transitions), eaaf5ad3 (Multiple Artificial Lights, Wet Surfaces, Rainy Conditions).
  • Challenging Waymo (6 scenes): segment-10 (Night Scene, Artificial Lighting, Lens Flare), segment-30 (Night Scene, Rainy Conditions, Specular Reflections), segment-539 (High Dynamic Range, Low-Angle Sun, Sun Glare), segment-550 (High Dynamic Range, Storefront Reflections, Hard Shadows), segment-561 (Low-Angle Sun, Lens Flare, Overexposure), segment-570 (Rainy Conditions, Low Light, Motion Blur).
  • Challenging Pandaset (6 scenes): 19, 21, 29, 48, 52 (all Low-Angle Sun with Hard/Deep Shadows or High Dynamic Range), 63 (Night Scene, Multiple Light Sources, Glare).
C.4 Vehicle-Specific Metrics
To specifically evaluate performance on challenging, highly reflective surfaces such as car windows, we used semantic masks to isolate vehicles. Table 19 shows that our method improves not only appearance (PSNR) but also, crucially, geometry (CD) for these specific objects. This supports our claim that our method mitigates the negative geometric impact of such severe view-dependent effects.